Jim Carrey and his pet budgie go everywhere together. One afternoon Jim Carrey decides to go for a walk. Jim Carrey walks 6.0 x 102 m [N] in 4.0 minutes, then walks 0.20 km [E] in 2.5 minutes.

Determine Jim Carrey's average velocity


1
Expert's answer
2021-08-02T07:33:03-0400

Average velocity is total displacement over total time. The displacement:


"D=\sqrt{600^2+200^2}=632\text{ m}."


The time:


"t=4+2.5=6.5\text{ min}."

Average velocity:


"v_\text{av}=\frac{632}{6.5}=97.2\text{ m/min}."